# Context-measure: Contextualizing Metric for Camouflage

### v1

@@ 1 Introduction
Evaluation metrics are fundamental to image segmentation community, serving both as a basis for model benchmarking, e.g., MS COCO [lin2014microsoft], PASCAL VOC [everingham2010pascal] and a catalyst for methodological innovation [zhao2019optimizing, milletari2016v, sudre2017generalised]. As a field evolves, it naturally calls for specialized evaluation metrics that align with its distinctive problem settings and evaluation philosophy. This study takes camouflaged object segmentation (COS) [fan2020camouflaged, fan2021concealed, ji2023deep, zhou2025camsam, ji2023sam] as a representative case.
Unlike generic segmentation, COS requires models to segment main targets that are blended with their context, exhibiting low target-background contrast and high structural homogeneity. However, existing evaluation practice in COS still relies on generic segmentation metrics, originally tailored for salient scenarios [borji2015salient, fan2022salient], such as the $F_{\beta}$ [alpert2011image, arbelaez2010contour], $S_{\alpha}$ [fan2021structure, fan2017structure], and $E_{\phi}$ [fan2021cognitive, fan2018enhanced]. These saliency-era metrics assess model performance solely based on the comparison between the model-predicted foreground maps (FM) and the manually annotated ground-truth (GT), without accounting for the visual complexity introduced by the context in which the targets are embedded. This naturally raises a question: “Should saliency and camouflage share the same metrics?” Before presenting the answer, please refer to Fig. 1: two morphologically identical butterflies exhibit completely different visual characteristics – the red butterfly is camouflaged, while the blue butterfly is salient. The primary reason for this discrepancy lies in the different contexts in which they are embedded. Based on this example, we can conclude that camouflage is inherently context-dependent, whereas these saliency-era metrics are context-independent. At this point, the answer to the question becomes clear: camouflage requires context-dependent evaluation metrics, distinct from those designed for saliency.
@@
To evaluate the cognitive alignment of metric prediction with human perception, we curate a new dataset containing 600 human-ranked predictions. Compared with existing metrics, our measure achieves a 41% relative improvement in consistency with human judgment. Additionally, we conduct extensive experiments on three commonly used COS benchmarks. Results demonstrate that our Context-measure achieves superior results across four meta-measures [margolin2014evaluate].
@@ 2.1 Revisiting Saliency-era Metrics
Pixel-aware Metrics. These assess per-pixel accuracy by directly comparing the predicted FM with the GT. Mean absolute error (MAE) measures the average absolute FM-GT difference at the pixel level. Intersection over Union (IoU) quantifies the overlap between FM and GT, reflecting how accurately the model captures object shape and spatial extent. The F-measure ($F_{\beta}$) [alpert2011image, arbelaez2010contour], defined as the harmonic mean of precision and recall, balances these two aspects to evaluate overall performance. When $\beta=1$, it reduces to the F1 score, which is mathematically related to the Jaccard Index (JI) [jaccard1901etude]. Notably, they satisfy the relation of $\text{IoU}=\text{JI}=\frac{\text{F1}}{2-\text{F1}}$, thus yielding identical rankings in practice and often being used interchangeably. The multiscale IoU (mIoU) [ahmadzadeh2021multiscale] extends IoU by incorporating multiple resolutions, enabling more comprehensive evaluation across scales and improving sensitivity to fine boundary details.
Region-aware Metrics. Unlike pixel-aware metrics, which assume all pixels are equally important, region-aware metrics address two perceptual flaws: E-Flaw & D-Flaw as in Table 1. The weighted F-measure ($F_{\beta}^{\omega}$) [margolin2014evaluate] mitigates these issues through distance-based weighting and error dependency modeling. The Structure-measure ($S_{\alpha}$) [fan2021structure, fan2017structure] integrates both region- and object-aware components, providing a more holistic evaluation of structural integrity in FMs.
Image-aware Metrics. These utilize global statistics to evaluate the holistic FM-GT alignment. For instance, the enhanced-alignment measure ($E_{\phi}$) [fan2018enhanced, fan2021cognitive] combines local pixel-level alignment with global image-level statistics, yielding results that are more consistent with human visual perception of foreground segmentation quality.
Table 1: Comparison of metric properties for foreground evaluation. Asm means whether the prediction and ground-truth are evaluated asymmetrically. E-Flaw (equal-importance flaw [margolin2014evaluate]) refers to the metric penalizes all erroneous pixels equally, regardless of perceptual significance. D-Flaw (dependence flaw [margolin2014evaluate]) assumes pixel-wise independence, thus ignoring spatial correlations. Image/Region/Pixel specify the perceptual hierarchy where a metric operates. Cam denotes the camouflaged scenarios for which the metric was originally designed.
@@
IoU/JI [jaccard1901etude]
$F_{\beta}$ [alpert2011image, arbelaez2010contour]
mIoU [ahmadzadeh2021multiscale]
$F_{\beta}^{\omega}$ [margolin2014evaluate]
$S_{\alpha}$ [fan2017structure, fan2021structure]
$E_{\phi}$ [fan2018enhanced, fan2021cognitive]
@@ 3.3 Contextualizing Context-measure for COS
Motivation. Existing camouflage quantification methods [lamdouar2023making, das2025camouflage] typically operate at the object level, measuring how well an entire object blends into its surroundings. However, in COS tasks, camouflage is inherently spatially non-uniform, different parts of the same object may exhibit varying camouflage degrees depending on their local color contrast with the background. This motivates us a pixel-aware quantification in the LAB color space, whose perceptual uniformity better aligns with human visual sensitivity. Compared to RGB, the LAB space decouples luminance and chromaticity, allowing us faithfully estimate color difference [sharma2005ciede2000] between object regions and their surroundings. Next, we detail two consecutive stages for our new method.
@@
Stage I: Contextual Overpainting. We proceed through this stage in four steps. (a) Extraction. Different from Lamdouar et al. [lamdouar2023making], we begin by expanding the object region outward by $k$ pixels through morphological dilation, thereby defining the expanded band as its contextual surroundings. From both the object and its surrounding context, we densely extract all overlapping $N\times N$ patches. (b) Embedding. Each patch is then embedded as a feature vector, obtained by concatenating its flattened LAB color values with standardized spatial coordinates, the latter scaled by a factor $\lambda$ to balance spatial and chromatic contributions. (c) Search. For every object feature vector, we perform an approximate nearest neighbor algorithm (ANN) to search the most compatible feature vector within the contextual surroundings. (d) Overpainting. Finally, the matched feature vector is projected back to the image space and used to overpaint the corresponding object region, resulting in a contextually harmonized reconstruction. This stage is illustrated in Fig. 5.
@@
where $\gamma$ modulates the nonlinearity. Here, $\Delta$ denotes the CIEDE2000 color difference [sharma2005ciede2000] between the repainted object $R_{lab}$ and the original object $I_{lab}$ in the LAB space. The visualized heatmaps of camouflage degree $\mathcal{D}$ are presented in Fig. 4, where higher values correspond to stronger camouflage, and lower values denote weaker camouflage. We further contextualize metric for camouflage by weighting $R(X|Y)$ (Eqn. 7) with camouflage degree $\mathcal{D}$. Accordingly, the reverse process can be reformulated as:
@@ 4.1 Meta-Measure 1: Human Ranking
$\bullet$ Evaluation Protocol. We compute the Spearman rank correlation coefficient $\rho\in[-1,1]$ [best1975algorithm] to quantify the consistency. For easier interpretation, we follow [margolin2014evaluate] in defining $\theta=1-\rho\in[0,2]$, where lower values indicates better ranking consistency. In addition, to facilitate comparative evaluation, this experiment is conducted on two datasets: CamoHR and FMDatabase [fan2018enhanced] – the latter being a binary foreground map dataset in salient scenarios that includes human-annotated rankings.
Table 2: Quantitative analysis of evaluation measures across 4 meta-measures. Only IoU results are reported, as IoU, F1, and JI are equivalent according to §2. Additionally, since the FMDatabase does not provide original images, $C_{\beta}^{\omega}$ cannot be computed on this dataset. The best and second-best results are highlighted in red and blue, respectively. $M$: Meta-Measure; DA#1: COD10K [fan2020camouflaged]; DA#2: NC4K [lv2021simultaneously]; DA#3: Trans10K [xie2020segmenting], $\text{Ours}^{1}$: $C_{\beta}$; $\text{Ours}^{2}$: $C_{\beta}^{\omega}$; 0.01%  : $\leq$0.01%.
@@ 5 Future Prospects
$\bullet$ Contextualizing for more scenarios. Beyond camouflage, our measure can be applied to broader domains. As shown in Fig. 9, without modifying any parameters, we observed that the heatmap generated by our quantification method performs well across various real scenarios [wang2024depth, fan2023advances, ji2025colonx]. Therefore, we believe our measure is adaptable to diverse applications by tailoring the spatial quantification based on the unique characteristics of task-specific data.
@@ Appendix A Revisiting Saliency-era Metrics
$\bullet$ F-measure ($F_{\beta}$). The F-measure [alpert2011image, arbelaez2010contour] is a standard evaluation metric that balances Precision and Recall when comparing the predicted mask with the ground-truth mask. Its computation depends on four commonly used pixel-level quantities: true positives ($TP$) — foreground pixels correctly predicted as foreground; true negatives ($TN$) — background pixels correctly predicted as background; false positives ($FP$) — background pixels incorrectly predicted as foreground; and false negatives ($FN$) — foreground pixels missed by the prediction. These quantities are computed as:
@@
$\bullet$ Weighted F-measure ($F_{\beta}^{\omega}$). The weighted F-measure, proposed by Margolin et al.  [margolin2014evaluate], addresses the inherent limitations of the F-measure and provides a more faithful quantification of foreground maps. Let the error vector be defined as $E=|Y^{\prime}-X^{\prime}|$. The weighted error vector is computed as:
@@
$\bullet$ Structure measure ($S_{\alpha}$). The $S_{\alpha}$ [fan2017structure] evaluates the structural consistency between the predicted mask and the ground-truth mask by jointly accounting for object-level integrity and region-level spatial layout. To compute the region-aware component, the image is partitioned into $K$ subregions, and a structural similarity score is computed in each region following an SSIM-inspired formulation. Let $\bar{X}$ and $\bar{Y}$ denote the mean values of $X$ and $Y$, and let $\sigma_{X}$, $\sigma_{Y}$, and $\sigma_{XY}$ be the standard deviations and covariance of $X^{\prime}$ and $Y^{\prime}$. The structural similarity for a given region is defined as:
@@
$\bullet$ Enhanced-alignment measure ($E_{\phi}$). The E-measure [fan2018enhanced] is a widely used saliency-evaluation metric designed to jointly capture both pixel-level matching and image-level statistics. As a first step, it constructs the bias matrices $\phi_{X}$ and $\phi_{Y}$ for $X$ and $Y$, by subtracting their mean values to remove global intensity bias:

### v2

@@ 1 Introduction
Evaluation metrics are fundamental to image segmentation community, serving both as a basis for model benchmarking, e.g., MS COCO [25], PASCAL VOC [7] and a catalyst for methodological innovation [35, 28, 32]. As a field evolves, it naturally calls for specialized evaluation metrics that align with its distinctive problem settings and evaluation philosophy. This study takes camouflaged object segmentation (COS) [13, 11, 21, 36, 22] as a representative case.
Unlike generic segmentation, COS requires models to segment main targets that are blended with their context, exhibiting low target-background contrast and high structural homogeneity. However, existing evaluation practice in COS still relies on generic segmentation metrics, originally tailored for salient scenarios [5, 15], such as the $F_{\beta}$ [2, 3], $S_{\alpha}$ [9, 8], and $E_{\phi}$ [12, 10]. These saliency-era metrics assess model performance solely based on the comparison between the model-predicted foreground maps (FM) and the manually annotated ground-truth (GT), without accounting for the visual complexity introduced by the context in which the targets are embedded. This naturally raises a question: “Should saliency and camouflage share the same metrics?” Before presenting the answer, please refer to Fig. 1: two morphologically identical butterflies exhibit completely different visual characteristics – the red butterfly is camouflaged, while the blue butterfly is salient. The primary reason for this discrepancy lies in the different contexts in which they are embedded. Based on this example, we can conclude that camouflage is inherently context-dependent, whereas these saliency-era metrics are context-independent. At this point, the answer to the question becomes clear: camouflage requires context-dependent evaluation metrics, distinct from those designed for saliency.
@@
To evaluate the cognitive alignment of metric prediction with human perception, we curate a new dataset containing 600 human-ranked predictions. Compared with existing metrics, our measure achieves a 41% relative improvement in consistency with human judgment. Additionally, we conduct extensive experiments on three commonly used COS benchmarks. Results demonstrate that our Context-measure achieves superior results across four meta-measures [27].
@@ 2.1 Revisiting Saliency-era Metrics
Pixel-aware Metrics. These assess per-pixel accuracy by directly comparing the predicted FM with the GT. Mean absolute error (MAE) measures the average absolute FM-GT difference at the pixel level. Intersection over Union (IoU) quantifies the overlap between FM and GT, reflecting how accurately the model captures object shape and spatial extent. The F-measure ($F_{\beta}$) [2, 3], defined as the harmonic mean of precision and recall, balances these two aspects to evaluate overall performance. When $\beta=1$, it reduces to the F1 score, which is mathematically related to the Jaccard Index (JI) [20]. Notably, they satisfy the relation of $\text{IoU}=\text{JI}=\frac{\text{F1}}{2-\text{F1}}$, thus yielding identical rankings in practice and often being used interchangeably. The multiscale IoU (mIoU) [1] extends IoU by incorporating multiple resolutions, enabling more comprehensive evaluation across scales and improving sensitivity to fine boundary details.
Region-aware Metrics. Unlike pixel-aware metrics, which assume all pixels are equally important, region-aware metrics address two perceptual flaws: E-Flaw & D-Flaw as in Table 1. The weighted F-measure ($F_{\beta}^{\omega}$) [27] mitigates these issues through distance-based weighting and error dependency modeling. The Structure-measure ($S_{\alpha}$) [9, 8] integrates both region- and object-aware components, providing a more holistic evaluation of structural integrity in FMs.
Image-aware Metrics. These utilize global statistics to evaluate the holistic FM-GT alignment. For instance, the enhanced-alignment measure ($E_{\phi}$) [10, 12] combines local pixel-level alignment with global image-level statistics, yielding results that are more consistent with human visual perception of foreground segmentation quality.
Table 1: Comparison of metric properties for foreground evaluation. Asm means whether the prediction and ground-truth are evaluated asymmetrically. E-Flaw (equal-importance flaw [27]) refers to the metric penalizes all erroneous pixels equally, regardless of perceptual significance. D-Flaw (dependence flaw [27]) assumes pixel-wise independence, thus ignoring spatial correlations. Image/Region/Pixel specify the perceptual hierarchy where a metric operates. Cam denotes the camouflaged scenarios for which the metric was originally designed.
@@
IoU/JI [20]
$F_{\beta}$ [2, 3]
mIoU [1]
$F_{\beta}^{\omega}$ [27]
$S_{\alpha}$ [8, 9]
$E_{\phi}$ [10, 12]
@@ 3.3 Contextualizing Context-measure for COS
Motivation. Existing camouflage quantification methods [24, 6] typically operate at the object level, measuring how well an entire object blends into its surroundings. However, in COS tasks, camouflage is inherently spatially non-uniform, different parts of the same object may exhibit varying camouflage degrees depending on their local color contrast with the background. This motivates us a pixel-aware quantification in the LAB color space, whose perceptual uniformity better aligns with human visual sensitivity. Compared to RGB, the LAB space decouples luminance and chromaticity, allowing us faithfully estimate color difference [31] between object regions and their surroundings. Next, we detail two consecutive stages for our new method.
@@
Stage I: Contextual Overpainting. We proceed through this stage in four steps. (a) Extraction. Different from Lamdouar et al. [24], we begin by expanding the object region outward by $k$ pixels through morphological dilation, thereby defining the expanded band as its contextual surroundings. From both the object and its surrounding context, we densely extract all overlapping $N\times N$ patches. (b) Embedding. Each patch is then embedded as a feature vector, obtained by concatenating its flattened LAB color values with standardized spatial coordinates, the latter scaled by a factor $\lambda$ to balance spatial and chromatic contributions. (c) Search. For every object feature vector, we perform an approximate nearest neighbor algorithm (ANN) to search the most compatible feature vector within the contextual surroundings. (d) Overpainting. Finally, the matched feature vector is projected back to the image space and used to overpaint the corresponding object region, resulting in a contextually harmonized reconstruction. This stage is illustrated in Fig. 5.
@@
where $\gamma$ modulates the nonlinearity. Here, $\Delta$ denotes the CIEDE2000 color difference [31] between the repainted object $R_{lab}$ and the original object $I_{lab}$ in the LAB space. The visualized heatmaps of camouflage degree $\mathcal{D}$ are presented in Fig. 4, where higher values correspond to stronger camouflage, and lower values denote weaker camouflage. We further contextualize metric for camouflage by weighting $R(X|Y)$ (Eqn. 7) with camouflage degree $\mathcal{D}$. Accordingly, the reverse process can be reformulated as:
@@ 4.1 Meta-Measure 1: Human Ranking
$\bullet$ Evaluation Protocol. We compute the Spearman rank correlation coefficient $\rho\in[-1,1]$ [4] to quantify the consistency. For easier interpretation, we follow [27] in defining $\theta=1-\rho\in[0,2]$, where lower values indicates better ranking consistency. In addition, to facilitate comparative evaluation, this experiment is conducted on two datasets: CamoHR and FMDatabase [10] – the latter being a binary foreground map dataset in salient scenarios that includes human-annotated rankings.
Table 2: Quantitative analysis of evaluation measures across 4 meta-measures. Only IoU results are reported, as IoU, F1, and JI are equivalent according to §2. Additionally, since the FMDatabase does not provide original images, $C_{\beta}^{\omega}$ cannot be computed on this dataset. The best and second-best results are highlighted in red and blue, respectively. $M$: Meta-Measure; DA#1: COD10K [13]; DA#2: NC4K [26]; DA#3: Trans10K [34], $\text{Ours}^{1}$: $C_{\beta}$; $\text{Ours}^{2}$: $C_{\beta}^{\omega}$; 0.01%  : $\leq$0.01%.
@@ 5 Future Prospects
$\bullet$ Contextualizing for more scenarios. Beyond camouflage, our measure can be applied to broader domains. As shown in Fig. 9, without modifying any parameters, we observed that the heatmap generated by our quantification method performs well across various real scenarios [33, 14, 23]. Therefore, we believe our measure is adaptable to diverse applications by tailoring the spatial quantification based on the unique characteristics of task-specific data.
@@ Appendix A Revisiting Saliency-era Metrics
$\bullet$ F-measure ($F_{\beta}$). The F-measure [2, 3] is a standard evaluation metric that balances Precision and Recall when comparing the predicted mask with the ground-truth mask. Its computation depends on four commonly used pixel-level quantities: true positives ($TP$) — foreground pixels correctly predicted as foreground; true negatives ($TN$) — background pixels correctly predicted as background; false positives ($FP$) — background pixels incorrectly predicted as foreground; and false negatives ($FN$) — foreground pixels missed by the prediction. These quantities are computed as:
@@
$\bullet$ Weighted F-measure ($F_{\beta}^{\omega}$). The weighted F-measure, proposed by Margolin et al.  [27], addresses the inherent limitations of the F-measure and provides a more faithful quantification of foreground maps. Let the error vector be defined as $E=|Y^{\prime}-X^{\prime}|$. The weighted error vector is computed as:
@@
$\bullet$ Structure measure ($S_{\alpha}$). The $S_{\alpha}$ [8] evaluates the structural consistency between the predicted mask and the ground-truth mask by jointly accounting for object-level integrity and region-level spatial layout. To compute the region-aware component, the image is partitioned into $K$ subregions, and a structural similarity score is computed in each region following an SSIM-inspired formulation. Let $\bar{X}$ and $\bar{Y}$ denote the mean values of $X$ and $Y$, and let $\sigma_{X}$, $\sigma_{Y}$, and $\sigma_{XY}$ be the standard deviations and covariance of $X^{\prime}$ and $Y^{\prime}$. The structural similarity for a given region is defined as:
@@
$\bullet$ Enhanced-alignment measure ($E_{\phi}$). The E-measure [10] is a widely used saliency-evaluation metric designed to jointly capture both pixel-level matching and image-level statistics. As a first step, it constructs the bias matrices $\phi_{X}$ and $\phi_{Y}$ for $X$ and $Y$, by subtracting their mean values to remove global intensity bias: